OpenStudy (anonymous):

A conical tent made of canvas has a base that is 34 feet across and a slant height of 12 feet. To the nearest whole unit, what is the area of the canvas, including the floor? Use 3.14 for π

OpenStudy (austinl):

Ok, so we have area formula, \[A = \pi r^2 + \pi r \sqrt{r^2 + h^2}\]  We need to solve for "h", we can do so using the Pythagorean Theorem

OpenStudy (austinl):

Does this make sense?

OpenStudy (anonymous):

yeahhhhh thank you !!!!!!!!!

OpenStudy (austinl):

no problem!!

OpenStudy (amistre64):